(3) A 20 cm long steel tube of 15 cm internal diameter and 1.0cm thickness is surrounded by a brass tube of the same length and thickness. The tubes carry an axial load of 150 kN. Estimate the load carried by each.
Es = 21 x 106 N/cm2 Eb = 10 x 106 N/cm2.
1
Expert's answer
2020-08-17T09:20:40-0400

Given:

Load P: 150 kN

Length L: 20 cm

Steel tube

Internal diameter "d_s": 15 cm

External diameter "D_s": 17 cm

Es = 21x106 N/cm2

Brass tube

Internal diameter "d_b": 17 cm

External diameter "D_b": 19 cm

Eb = 10x106 N/cm2


solution

Area of steel and brass tubes

"A_{s}=\frac{\pi}{4}(D^{2}_{s}-d^{2}_{s}=50.27cm^{2}"

"A_{b}=\frac{\pi}{4}(D^{2}_{b}-d^{2}_{b})=56.55cm^{2}"

Under the same load strain "\epsilon" in steel equals the strain in brass:

"\frac{\sigma_{s}}{E_{s}}=\frac{\sigma_{b}}{E_{b}}"

"\sigma_{s}=\sigma_{b}\frac{E_{s}}{E_{b}}"

Load on steel added to the load on brass equals the total load:

"P_{s}+P_{b}=P"

"\sigma_{s}A_{s}+\sigma_{b}A_{b}=P"

"\sigma_{b}(\frac{E_{s}}{E_{b}}A_{s}+A_{b})=P"

Load carried by brass and steel respectively:

"P_{b}=\sigma_{s}A_{b}=\frac{PA_{b}}{\frac{E_{s}}{E_{b}}A_{s}+A_{b}}=\frac{PE_{b}A_{b}}{E_{s}A_{s}+E_{b}A_{b}}=52.3kN"

"P_{s}=\sigma_{s}A_{s}=\frac{PA_{s}}{\frac{E_{s}}{E_{b}}A_{s}+A_{b}}=\frac{PE_{s}A_{s}}{E_{s}A_{s}+E_{b}A_{b}}=97.7kN"